# A Note on Restricted Selection Set from Random Interval

### v1

@@ 2.4. Auxiliary rearrangement lemmas
Lemma 2.5 (Conditional rearrangement / quantile bound[LiebLoss2001]).
@@
Lemma 2.6 (Quantile–area identity [HardyLittlewoodPolya1952]).
@@
Set $Q(u):=F_{Z}^{-1}(u)$ for $u\in(0,1)$ and fix $\alpha\in(0,1)$. Since $Z\geq 0$ almost surely, we have $Q(u)\geq 0$ for all $u\in(0,1)$. Using the layer-cake representation and Tonelli’s theorem,
$\int_{0}^{\alpha}Q(u)\,du$ $=\int_{0}^{\alpha}\left(\int_{0}^{Q(u)}dt\right)du$
$=\int_{0}^{\infty}\left(\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du\right)dt.$
Thus, it suffices to compute, for each fixed $t\geq 0$, the inner integral
$\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du$
as a function of $t$.
By definition of the quantile function,
$Q(u)\leq t\quad\Longleftrightarrow\quad u\leq F_{Z}(t),$
because $F_{Z}(t)=\sup\{u\in[0,1]:Q(u)\leq t\}$ for each $t$. Hence
$t\leq Q(u)\quad\Longleftrightarrow\quad u\geq F_{Z}(t).$
Therefore, for fixed $t\geq 0$,
$\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du=\lambda\bigl(\{u\in[0,\alpha]:u\geq F_{Z}(t)\}\bigr),$
where $\lambda$ denotes Lebesgue measure on $\mathbb{R}$. This set is empty if $F_{Z}(t)\geq\alpha$, and it is the interval $F_{Z}(t),\alpha]$ if $F_{Z}(t)<\alpha$. Consequently,
$\int_{0}^{\alpha}\mathbf{1}\{t\leq Q(u)\}\,du=\begin{cases}\alpha-F_{Z}(t),&\text{if }F_{Z}(t)<\alpha,\\[3.0pt] 0,&\text{if }F_{Z}(t)\geq\alpha,\end{cases}$
which is exactly $(\alpha-F_{Z}(t))_{+}$. Plugging this back into the layer-cake formula gives
@@
which is (2.8). ∎
2.5. Identification of $\mathbb{E}(y^{*})$ under a median restriction
We now consider the effect of knowing a median on the range of admissible expectations. Let $Y=[y_{L},y_{U}]$ be a random interval with $\mathbb{E}|y_{L}|+\mathbb{E}|y_{U}|<\infty$, and suppose we know that some selection $y\in\operatorname{\mathbf{Sel}}(Y)$ has population median $m\in\mathbb{R}$ in the sense that
@@ 2.5. Identification of $\mathbb{E}(y^{*})$ under a median restriction
Assume feasibility, $p_{-}\leq\tfrac{1}{2}$ and $p_{+}\leq\tfrac{1}{2}$. Then the identified set of means under the median restriction
$\mathcal{S}_{m}:=\{\,y\in\operatorname{\mathbf{Sel}}(Y):\ \operatorname*{Med}(y)=m\,\}$
@@
$\mathbb{E}[Y\mid\mathcal{S}_{m}]\;=\;\Big[\,\mathbb{E}_{\min}(m),\ \mathbb{E}_{\max}(m)\,\Big],$ (2.9)
@@
$\bar{y}(\omega):=\begin{cases}m(\omega),&\omega\in S_{-},\\ y_{U}(\omega),&\omega\in M\setminus S_{-},\\ y_{U}(\omega),&\omega\in A_{-}\cup A_{+}.\end{cases}$
Then $\bar{y}\in\operatorname{\mathbf{Sel}}(Y)$, $\mathbf{P}(\bar{y}\leq m)=p_{-}+\alpha_{-}(m)=\tfrac{1}{2}$, $\mathbf{P}(\bar{y}\geq m)=\tfrac{1}{2}$, and by construction
@@
$\underline{y}(\omega):=\begin{cases}m(\omega),&\omega\in S_{+},\\ y_{L}(\omega),&\omega\in M\setminus S_{+},\\ y_{L}(\omega),&\omega\in A_{-}\cup A_{+}.\end{cases}$
Then $\underline{y}\in\operatorname{\mathbf{Sel}}(Y)$, $\mathbf{P}(\underline{y}\geq m)=p_{+}+\alpha_{+}(m)=\tfrac{1}{2}$, $\mathbf{P}(\underline{y}\leq m)=\tfrac{1}{2}$, and
@@
We interpret $m$ as the known population median of a measurable selection $y\in\operatorname{\mathbf{Sel}}(Y)$, in the sense that $\mathbf{P}(y\leq m)\geq\tfrac{1}{2}$ and $\mathbf{P}(y\geq m)\geq\tfrac{1}{2}$.
In this configuration the identification region for the mean simplifies to
$\mathbb{E}(y^{*})\in\big[\,\mathbb{E}(y_{L})+S_{L},\ \mathbb{E}(y_{U})-S_{U}\,\big],$
@@
Remark 2.10 (Derivation of $S_{L}$ and $S_{U}$).
For completeness, we show how the expressions for $S_{L}$ and $S_{U}$ in Example 2.9 arise from the general formulas in Proposition 2.7.
A useful identity is that for any nonnegative real-valued random variable $Z$ with CDF $F_{Z}$ and any $\alpha\in(0,1)$ by Lemma 2.6
where $(x)_{+}:=\max\{x,0\}$. This follows from a standard layer-cake (rearrangement) argument.
In the notation of Proposition 2.7, take
$L_{m}:=m-y_{L},\qquad U_{m}:=y_{U}-m,$
so that $L_{m},U_{m}\geq 0$ and, writing $F_{L},F_{U}$ for the CDFs of $y_{L},y_{U}$,
$F_{L_{m}}(\ell)=\mathbf{P}(L_{m}\leq\ell)=\mathbf{P}(y_{L}\geq m-\ell)=1-F_{L}(m-\ell),$
$F_{U_{m}}(\ell)=\mathbf{P}(U_{m}\leq\ell)=\mathbf{P}(y_{U}\leq m+\ell)=F_{U}(m+\ell).$
In the schematic configuration we have $m_{L}<m<m_{U}$, so the median shortfalls are symmetric and Proposition 2.7 reduces to
$\mathbb{E}_{\min}(m)=\mathbb{E}[y_{L}]+\int_{0}^{1/2}F^{-1}_{L_{m}}(u)\,du,\qquad\mathbb{E}_{\max}(m)=\mathbb{E}[y_{U}]-\int_{0}^{1/2}F^{-1}_{U_{m}}(u)\,du.$
Apply (2.12) with $Z=L_{m}$ and $\alpha=\tfrac{1}{2}$:
$\int_{0}^{1/2}F^{-1}_{L_{m}}(u)\,du=\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{L_{m}}(\ell)\bigr)_{+}\,d\ell=\int_{0}^{\infty}\bigl(F_{L}(m-\ell)-\tfrac{1}{2}\bigr)_{+}\,d\ell.$
With the change of variable $t=m$-$\ell$ and $d\ell=-dt$, this becomes
$\int_{0}^{1/2}F^{-1}_{L_{m}}(u)\,du=\int_{-\infty}^{m}\bigl(F_{L}(t)-\tfrac{1}{2}\bigr)_{+}\,dt=\int_{m_{L}}^{m}\bigl(F_{L}(t)-\tfrac{1}{2}\bigr)\,dt,$
since $F_{L}(t)\leq\tfrac{1}{2}$ for $t\leq m_{L}$ and the integrand is zero there. This yields
$\mathbb{E}_{\min}(m)=\mathbb{E}[y_{L}]+\int_{m_{L}}^{m}\bigl(F_{L}(t)-\tfrac{1}{2}\bigr)\,dt=\mathbb{E}[y_{L}]+S_{L}.$
Similarly, with $Z=U_{m}$ and $\alpha=\tfrac{1}{2}$,
$\int_{0}^{1/2}F^{-1}_{U_{m}}(u)\,du=\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{U_{m}}(\ell)\bigr)_{+}\,d\ell=\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{U}(m+\ell)\bigr)_{+}\,d\ell.$
Setting $t=m+\ell$ gives
$\int_{0}^{1/2}F^{-1}_{U_{m}}(u)\,du=\int_{m}^{\infty}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)_{+}\,dt=\int_{m}^{m_{U}}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)\,dt,$
because $F_{U}(t)\geq\tfrac{1}{2}$ for $t\geq m_{U}$. Hence
$\mathbb{E}_{\max}(m)=\mathbb{E}[y_{U}]-\int_{m}^{m_{U}}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)\,dt=\mathbb{E}[y_{U}]-S_{U}.$
This shows how the “cost” terms $S_{L}$ and $S_{U}$ in Example 2.9 arise from the general quantile-integral representation in Proposition 2.7 by a simple change of variables.
@@
2.6. Identification of $\operatorname*{Med}(y^{*})$ under a mean restriction
We now consider the reverse question: given a mean restriction $\mathbb{E}(y^{*})=\kappa$, does this necessarily sharpen the benchmark median interval $\Theta_{M}^{*}=[\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$? The next example shows that the answer is negative in general.
Example 2.11 (Two–state random interval with symmetry around $\kappa$).
Fix $\kappa\in\mathbb{R}$ and $d>0$. On a probability space with two equiprobable states $\{\omega_{1},\omega_{2}\}$, define the random interval
$Y(\omega_{1})=[\,\kappa-2d,\ \kappa\,],\qquad Y(\omega_{2})=[\,\kappa,\ \kappa+2d\,],$
equivalently
$(y_{L},y_{U})=\begin{cases}(\kappa-2d,\ \kappa),&\omega=\omega_{1},\\ (\kappa,\ \kappa+2d),&\omega=\omega_{2},\end{cases}\qquad\mathbf{P}(\omega_{1})=\mathbf{P}(\omega_{2})=\tfrac{1}{2}.$
Then $y_{L}$ takes values $\{\kappa-2d,\kappa\}$ and $y_{U}$ takes values $\{\kappa,\kappa+2d\}$ (each with probability $1/2$), so
$\operatorname*{Med}(y_{L})=\kappa-2d,\qquad\operatorname*{Med}(y_{U})=\kappa,$
and the unrestricted median bound is $\operatorname*{Med}(y^{*})\in[\kappa-2d,\kappa]$.
Now fix any $m\in[\kappa-2d,\kappa]$ and define the selection
$y(\omega_{1}):=m,\qquad y(\omega_{2}):=2\kappa-m.$
Because $m\in[\kappa-2d,\kappa]$ and $2\kappa-m\in[\kappa,\kappa+2d]$, we have $y(\omega_{i})\in Y(\omega_{i})$ for $i=1,2$, hence $y\in\operatorname{\mathbf{Sel}}(Y)$. Moreover,
$\mathbb{E}[y]=\tfrac{1}{2}\,m+\tfrac{1}{2}\,(2\kappa-m)=\kappa,$
so $y$ satisfies the mean restriction. Since the distribution of $y$ places probability $1/2$ at $m$ and $1/2$ at $2\kappa-m$ with $m\leq\kappa\leq 2\kappa-m$, it follows that
$\mathbf{P}(y\leq m)=\tfrac{1}{2},\qquad\mathbf{P}(y\geq m)=1,$
hence $\operatorname*{Med}(y)=m$ (under the standard definition $\mathbf{P}(y\leq m)\geq\tfrac{1}{2}$ and $\mathbf{P}(y\geq m)\geq\tfrac{1}{2}$).
Because $m$ was arbitrary in $\kappa-2d,\kappa]$, we conclude
$\big\{\operatorname*{Med}(y):\,y\in\operatorname{\mathbf{Sel}}(Y),\ \mathbb{E}[y]=\kappa\big\}=[\kappa-2d,\ \kappa]=\big[\operatorname*{Med}(y_{L}),\,\operatorname*{Med}(y_{U})\big].$
Thus, in this random–interval design, knowing $\mathbb{E}(y^{*})=\kappa$ does not shrink the identified set for $\operatorname*{Med}(y^{*})$.

### v2

@@ 2.4. Auxiliary rearrangement lemmas
Lemma 2.5 (Conditional rearrangement / quantile bound[6]).
@@
Lemma 2.6 (Quantile–area identity [4]).
@@
Set $Q(u):=F_{Z}^{-1}(u)$ for $u\in(0,1)$ and fix $\alpha\in(0,1)$. Since $Z\geq 0$ almost surely, we have $Q(u)\geq 0$ for all $u\in(0,1)$. By the layer-cake representation and Tonelli’s theorem,
$\int_{0}^{\alpha}Q(u)\,du$ $=\int_{0}^{\alpha}\left(\int_{0}^{\infty}\mathbf{1}_{\{t<Q(u)\}}\,dt\right)du$
$=\int_{0}^{\infty}\left(\int_{0}^{\alpha}\mathbf{1}_{\{t<Q(u)\}}\,du\right)dt.$
For the left-continuous quantile function
$Q(u)=\inf\{x\in\mathbb{R}:F_{Z}(x)\geq u\},$
we have, for every $t\geq 0$ and $u\in(0,1)$,
$t<Q(u)\quad\Longleftrightarrow\quad F_{Z}(t)<u.$
Hence, for fixed $t\geq 0$,
$\int_{0}^{\alpha}\mathbf{1}_{\{t<Q(u)\}}\,du=\int_{0}^{\alpha}\mathbf{1}_{\{F_{Z}(t)<u\}}\,du=(\alpha-F_{Z}(t))_{+}.$
Substituting this into the previous display gives
@@
which is exactly (2.9). ∎
2.5. Bounds on $\mathbb{E}(y)$ over $\operatorname{\mathbf{Sel}}(Y\mid m)$
We now consider the effect of knowing a median on the range of admissible expectations. Let $Y=[y_{L},y_{U}]$ be a random interval with $\mathbb{E}|y_{L}|+\mathbb{E}|y_{U}|<\infty$, and consider the restricted selection set $\operatorname{\mathbf{Sel}}(Y\mid m)$ of all selections with median $m$ in the sense that
@@ 2.5. Bounds on $\mathbb{E}(y)$ over $\operatorname{\mathbf{Sel}}(Y\mid m)$
Assume feasibility, $p_{-}\leq\tfrac{1}{2}$ and $p_{+}\leq\tfrac{1}{2}$. The attainable set of means under the median restriction
$\operatorname{\mathbf{Sel}}(Y\mid m)=\{y\in\operatorname{\mathbf{Sel}}(Y):m\in\operatorname*{Med}(y)\}.$
@@
$\{\mathbb{E}[y]:y\in\mathcal{S}_{m}\}\;=\;\Big[\,\mathbb{E}_{\min}(m),\ \mathbb{E}_{\max}(m)\,\Big],$ (2.10)
@@
$\bar{y}(\omega):=\begin{cases}m,&\omega\in S_{-},\\ y_{U}(\omega),&\omega\in M\setminus S_{-},\\ y_{U}(\omega),&\omega\in A_{-}\cup A_{+}.\end{cases}$
Then $\bar{y}\in\operatorname{\mathbf{Sel}}(Y)$. We verify the median condition $m\in\operatorname*{Med}(\bar{y})$.
On $A_{-}$ we have $\bar{y}=y_{U}<m$ (since $A_{-}=\{y_{U}<m\}$); on $S_{-}\subset M$ we have $\bar{y}=m$; and on $M\setminus S_{-}$ we have $\bar{y}=y_{U}\geq m$ (with equality when $y_{U}=m$). On $A_{+}$ we have $\bar{y}=y_{U}\geq y_{L}>m$. Therefore
$\mathbf{P}(\bar{y}\leq m)\;=\;\mathbf{P}(A_{-})+\mathbf{P}(S_{-})+\mathbf{P}\bigl(\{y_{U}=m\}\cap(M\setminus S_{-})\bigr)\;\geq\;p_{-}+\alpha_{-}(m)\;=\;\tfrac{1}{2}.$
For the other half: on $A_{-}$ we have $\bar{y}=y_{U}<m$, so $\bar{y}\geq m$ fails there; everywhere else $\bar{y}\geq m$. Hence
$\mathbf{P}(\bar{y}\geq m)\;=\;1-\mathbf{P}(A_{-})\;=\;1-p_{-}\;\geq\;\tfrac{1}{2},$
using the feasibility assumption $p_{-}\leq\tfrac{1}{2}$. Thus $m\in\operatorname*{Med}(\bar{y})$, and by construction
@@
$\underline{y}(\omega):=\begin{cases}m&\omega\in S_{+},\\ y_{L}(\omega),&\omega\in M\setminus S_{+},\\ y_{L}(\omega),&\omega\in A_{-}\cup A_{+}.\end{cases}$
Then $\underline{y}\in\operatorname{\mathbf{Sel}}(Y)$. On $A_{+}$ we have $\underline{y}=y_{L}>m$; on $S_{+}\subset M$ we have $\underline{y}=m$; and on $M\setminus S_{+}$ we have $\underline{y}=y_{L}\leq m$ (with equality when $y_{L}=m$). On $A_{-}$ we have $\underline{y}=y_{L}\leq y_{U}<m$. Therefore
$\mathbf{P}(\underline{y}\geq m)\;=\;\mathbf{P}(A_{+})+\mathbf{P}(S_{+})+\mathbf{P}\bigl(\{y_{L}=m\}\cap(M\setminus S_{+})\bigr)\;\geq\;p_{+}+\alpha_{+}(m)\;=\;\tfrac{1}{2},$
and
$\mathbf{P}(\underline{y}\leq m)\;=\;1-\mathbf{P}(A_{+})\;=\;1-p_{+}\;\geq\;\tfrac{1}{2}.$
Thus $m\in\operatorname*{Med}(\underline{y})$, and
@@
We fix $m$ as the median constraint, so that $\operatorname{\mathbf{Sel}}(Y\mid m)$ consists of all $y\in\operatorname{\mathbf{Sel}}(Y)$ satisfying that $\mathbf{P}(y\leq m)\geq\tfrac{1}{2}$ and $\mathbf{P}(y\geq m)\geq\tfrac{1}{2}$.
In this configuration the bounded region for the mean simplifies to
$\mathbb{E}[y]\in\big[\,\mathbb{E}(y_{L})+S_{L},\ \mathbb{E}(y_{U})-S_{U}\,\big]\qquad\text{for all }y\in\operatorname{\mathbf{Sel}}(Y\mid m),$
@@
Remark 2.10 (Derivation of the bounds in Example 2.9).
We show how to pass from the general formulas of Proposition 2.7 to explicit CDF integrals. The key tool is the quantile–area identity (Lemma 2.6): for any nonnegative random variable $Z$ with CDF $F_{Z}$ and any $\alpha\in(0,1)$,
$\int_{0}^{\alpha}F_{Z}^{-1}(u)\,du\;=\;\int_{0}^{\infty}(\alpha-F_{Z}(t))_{+}\,dt.$ (2.13)
Write $p_{-}=\mathbf{P}(y_{U}<m)$, $p_{+}=\mathbf{P}(y_{L}>m)$, $p_{0}=\mathbf{P}(y_{L}\leq m\leq y_{U})=1-p_{-}-p_{+}$, and $\alpha_{-}=\tfrac{1}{2}-p_{-}$, $\alpha_{+}=\tfrac{1}{2}-p_{+}$ (assuming feasibility $p_{-},p_{+}\leq\tfrac{1}{2}$).
Upper endpoint. By Proposition 2.7,
$\mathbb{E}_{\max}(m)\;=\;\mathbb{E}[y_{U}]\;-\;p_{0}\!\int_{0}^{\,\alpha_{-}/p_{0}}\!F^{-1}_{U_{m}\mid M}(u)\,du,$
where $U_{m}=y_{U}-m$ and the quantile is conditional on $M=\{y_{L}\leq m\leq y_{U}\}$. Applying identity (2.13) on the conditional space $(M,\mathbf{P}_{M})$ with $Z=U_{m}$ and $\alpha=\alpha_{-}/p_{0}$ gives
$p_{0}\!\int_{0}^{\alpha_{-}/p_{0}}\!F^{-1}_{U_{m}\mid M}(u)\,du\;=\;p_{0}\!\int_{0}^{\infty}\Bigl(\frac{\alpha_{-}}{p_{0}}-F_{U_{m}\mid M}(\ell)\Bigr)_{+}d\ell.$
Since $F_{U_{m}\mid M}(\ell)=\mathbf{P}_{M}(y_{U}-m\leq\ell)=\mathbf{P}_{M}(y_{U}\leq m+\ell)=\frac{\mathbf{P}(y_{U}\leq m+\ell,\;y_{L}\leq m\leq y_{U})}{p_{0}}$, and writing $G(\ell):=\mathbf{P}(y_{U}\leq m+\ell,\;y_{L}\leq m\leq y_{U})$ so that $F_{U_{m}\mid M}(\ell)=G(\ell)/p_{0}$, we obtain
$p_{0}\!\int_{0}^{\alpha_{-}/p_{0}}\!F^{-1}_{U_{m}\mid M}(u)\,du\;=\;\int_{0}^{\infty}\bigl(\alpha_{-}-G(\ell)\bigr)_{+}\,d\ell.$
In the special case $p_{-}=p_{+}=0$ (equivalently $m\in[y_{L},y_{U}]$ a.s.), we have $p_{0}=1$, $\alpha_{-}=\alpha_{+}=\tfrac{1}{2}$, and $G(\ell)=F_{U}(m+\ell)$, so the integral reduces to
$\int_{0}^{\infty}\bigl(\tfrac{1}{2}-F_{U}(m+\ell)\bigr)_{+}\,d\ell\;=\;\int_{m}^{m_{U}}\bigl(\tfrac{1}{2}-F_{U}(t)\bigr)\,dt\;=:\;S_{U}.$
The analogous calculation for the lower endpoint yields $S_{L}=\int_{m_{L}}^{m}(F_{L}(t)-\tfrac{1}{2})\,dt$ under the same simplifying assumption.
In the chi-squared configuration of Example 2.9, $p_{-}$ and $p_{+}$ are strictly positive, so the simple formulas involving $S_{L}$ and $S_{U}$ are only approximate. However, the qualitative picture — the shaded areas between the CDFs and the 1/21/2–level being “costs” that shrink the Aumann range — remains valid. For exact numerical evaluation one should use the general conditional formulas of Proposition 2.7 directly.
@@
2.6. Attainable medians under a mean restriction
We now consider the reverse question: does the mean constraint $\mathbb{E}[y]=\kappa$ necessarily sharpen the benchmark median interval $\Theta_{M}=[\operatorname*{Med}(y_{L}),\operatorname*{Med}(y_{U})]$? The next example shows that the answer is negative in general.
Example 2.11 (Mean restriction need not shrink the attainable median set).
Fix $\kappa\in\mathbb{R}$ and $d>0$. Let $(\Omega,\mathcal{A},\mathbf{P})$ be a non-atomic probability space, and let $B\in\mathcal{A}$ satisfy $\mathbf{P}(B)=\tfrac{1}{2}$. Define the random interval
$Y(\omega)=\begin{cases}[\kappa-2d,\ \kappa],&\omega\in B,\\[5.69054pt] [\kappa,\ \kappa+2d],&\omega\in B^{c}.\end{cases}$
Equivalently,
$(y_{L},y_{U})=\begin{cases}(\kappa-2d,\ \kappa),&\omega\in B,\\ (\kappa,\ \kappa+2d),&\omega\in B^{c}.\end{cases}$
We first identify the benchmark median interval. Since
$y_{L}=\begin{cases}\kappa-2d,&\omega\in B,\\ \kappa,&\omega\in B^{c},\end{cases}\qquad y_{U}=\begin{cases}\kappa,&\omega\in B,\\ \kappa+2d,&\omega\in B^{c},\end{cases}$
each with probabilities $1/2$ and $1/2$, we have
$\operatorname*{Med}(y_{L})=[\kappa-2d,\kappa],\qquad\operatorname*{Med}(y_{U})=[\kappa,\kappa+2d].$
Hence the benchmark interval from (2.2) is
$\Theta_{M}=[\kappa-2d,\kappa+2d].$
Now fix any target median
$m\in[\kappa-2d,\kappa+2d].$
We show that there exists a selection $y\in\operatorname{\mathbf{Sel}}(Y)$ such that
$\mathbb{E}[y]=\kappa\qquad\text{and}\qquad m\in\operatorname*{Med}(y).$
Case 1: $m\in[\kappa-2d,\kappa]$. Define
$y(\omega)=\begin{cases}m,&\omega\in B,\\ 2\kappa-m,&\omega\in B^{c}.\end{cases}$
Since
$m\in[\kappa-2d,\kappa]\qquad\text{and}\qquad 2\kappa-m\in[\kappa,\kappa+2d],$
we have $y(\omega)\in Y(\omega)$ almost surely, so $y\in\operatorname{\mathbf{Sel}}(Y)$. Moreover,
$\mathbb{E}[y]=\tfrac{1}{2}\,m+\tfrac{1}{2}\,(2\kappa-m)=\kappa.$
Because $m\leq\kappa\leq 2\kappa-m$, the distribution of $y$ puts probability $1/2$ at $m$ and probability $1/2$ at $2\kappa-m$, hence
$\mathbf{P}(y\leq m)=\tfrac{1}{2},\qquad\mathbf{P}(y\geq m)=1.$
Therefore $m\in\operatorname*{Med}(y)$.
Case 2: $m\in[\kappa,\kappa+2d]$. Define
$y(\omega)=\begin{cases}2\kappa-m,&\omega\in B,\\ m,&\omega\in B^{c}.\end{cases}$
Now
$2\kappa-m\in[\kappa-2d,\kappa]\qquad\text{and}\qquad m\in[\kappa,\kappa+2d],$
so again $y\in\operatorname{\mathbf{Sel}}(Y)$. Also,
$\mathbb{E}[y]=\tfrac{1}{2}(2\kappa-m)+\tfrac{1}{2}m=\kappa.$
Since $2\kappa-m\leq\kappa\leq m$, we get
$\mathbf{P}(y\leq m)=1,\qquad\mathbf{P}(y\geq m)=\tfrac{1}{2},$
and therefore $m\in\operatorname*{Med}(y)$.
Since $m$ was arbitrary in $\kappa-2d,\kappa+2d]$, we conclude that
$\{\,m\in\mathbb{R}:\exists\,y\in\operatorname{\mathbf{Sel}}(Y)\text{ with }\mathbb{E}[y]=\kappa\text{ and }m\in\operatorname*{Med}(y)\,\}=[\kappa-2d,\kappa+2d]=\Theta_{M}.$
Thus, in this random-interval design, the mean restriction $\mathbb{E}[y]=\kappa$ does not shrink the benchmark attainable median interval.
In particular, the mean restriction need not reduce the set of attainable medians, even though it may reduce the set of attainable distributions.